Last fall, a gardener planted 62 iris bulbs. She found that only 48 of the bulbs bloomed in the spring.

Find the empirical probability that an iris bulb of this type will bloom. Give answer as a fraction in lowest terms.
How many of the bulbs should she plant next fall if she would like at least 53 to bloom?

I believe the first part is 48/62 but the 2nd not sure of

Answer by edjones(8007)

48/62=24/31
.
x/53=24/31 proportion
x=1272/31
=41.03 more than 41 should be planted.
.
Ed